If 0.231 mol of potassium fluoride (KF) is dissolved in 50.3 g of water, how much will the freezing point be lowered? (17.1°C)


1
Expert's answer
2021-04-28T02:46:06-0400

Solution:

The equation describing the change in freezing point from pure solvent to solution is:

ΔTf = i × Kf × m (for electrolytes)

where:

i = the number of dissolved particles (van’t Hoff factor).

m = molality = moles of solute per kilogram of solvent.

Kf is the molal freezing point depression constant of the solvent (1.86°C/m for water).


The van't Hoff factor is the number of particles (ions) formed by the compound in the solution.

The dissociation of KCl is given by,

KCl → K+ + Cl-

Therefore, van’t Hoff factor is 2 for KCl.


Molality of KF solution = Moles of KF / Kilograms of H2O

Molality of KF solution = (0.231 mol) / (0.0503 kg) = 4.5924 mol/kg = 4.5924 m


Thus,

ΔTf = (2) × (1.86°C/m) × (4.5924 m) = 17.08°C = 17.1°C

ΔTf = 17.1°C


Answer: The freezing point will be lowered by 17.1°C